Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability need help

Expert replies
by Winner2013 » Sun Aug 25, 2013 11:19 am
On a biased dice,any even number appears four times as frequently as any odd number. If the dice is rolled thrice what is the probability that the sum of the scores on those is more than 16?

1. 26/375 2. 112/375 3. 26/3375 4. 112/3375 5. 8/225. Answer is option 4
Join the discussion
Source: — Problem Solving |

by [email protected] » Sun Aug 25, 2013 12:30 pm
Hi Winner2013,

This question layers on the details, so you have to be very careful with how you organize it. There are a couple of different ways to do the math, but here's a way that allows you to see all of the possibilities:

First, let's deal with the fact that even numbers appear 4x as frequently as odd numbers.

Roll
1 x 1
2 x 4
3 x 1
4 x 4
5 x 1
6 x 4

Total possible outcomes per roll = 15

Now, we need to calculate the probability of what we WANT (a three-roll sum > 16): So, we need a total of 17 or 18.

To get a 17 or 18, we need one of the following possibilities:
665
656
566
666

Odds of 665: (4/15)(4/15)(1/15) = 4^2/15^3
Odds of 656: Same as above = 4^2/15^3
Odds of 566: Same as above = 4^2/15^3

Odds of 666: (4/15)(4/15)(4/15) = 4^3/15^3

Total odds: 16(3) + 64 = 112/15^3

Final Answer: 4

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Winner2013 » Sun Aug 25, 2013 1:05 pm
thank you Rich. I understood the logic better now. I have been following your posts and you help clarify concepts. thank you again. I would like to tell you what i thought about this problem just to understand how my thought process has been wrong, so that I don't repeat the same mistake again.

Chances of getting even are 4 times getting an odd.
So probability of getting even = 4/5
And p(odd) = 1/5

Now to get sum>16 it should be either 17 or 18
For 17- 3!/2! = 3 ways considering 3 arrangements
And 18 there is 1

So for 17 probability = (4/5 *4/5*1/5 ) *3
And for 18 = (4/5*4/5*4/5)

So total = (16*3 +64)/125 = 112/125

:-( i don't know where am i going wrong. Please help. Appreciate it. Thank you
Join the discussion

by [email protected] » Sun Aug 25, 2013 3:30 pm
Hi Winner2013,

The reason why your logic didn't quite work in this question is that you did not account for ALL of the numbers. To score 17 or 18 on three rolls, you CAN'T have any 1s, 2s, 3s or 4s. Your math doesn't account for that; it only accounts for the probability of rolling an odd (1/5) vs. rolling an even (4/5).

You DID correctly deduce that there are 3 ways to roll 17 and 1 way to roll 18. To adjust your calculation, you have to consider ALL the possible outcomes (so you divide each value by 15, not 5).

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Winner2013 » Sun Aug 25, 2013 5:18 pm
Yes you are right. Thank you very much Rich.
Join the discussion

by ganeshrkamath » Sun Aug 25, 2013 9:26 pm
Winner2013 wrote:On a biased dice,any even number appears four times as frequently as any odd number. If the dice is rolled thrice what is the probability that the sum of the scores on those is more than 16?

1. 26/375 2. 112/375 3. 26/3375 4. 112/3375 5. 8/225. Answer is option 4
Probability of even number appearing = 4p
Probability of odd number appearing = p
4p*3 + p*3 = 1
p = 1/15
4p = 4/15

To get more than 16 (i.e. 17 or 18) the numbers should be (6,6,5) or (6,6,6) in any order.
P(6,6,6) = 4/15*4/15*4/15 = 64/3375
P(6,6,5) = 4/15*4/15*1/15 = 16/3375 = P(6,5,6) = P(5,6,6)
Total probability = 64/3375 + 3*16/3375 = (64+48)/3375
= 112/3375

Choose [spoiler](4)[/spoiler]

Cheers
Every job is a self-portrait of the person who did it. Autograph your work with excellence.

Kelley School of Business (Class of 2016)
GMAT Score: 750 V40 Q51 AWA 5 IR 8
https://www.beatthegmat.com/first-attemp ... tml#688494
Join the discussion